OpenStudy (roberts.spurs19):

How would you find the line of intersection (in cartesian form) of the following planes please? 3i + j - 4k = 4 and -i + j - 2k = 2 Thank you!

OpenStudy (dan815):

hmm

OpenStudy (dan815):

i guess u cud mutliply the bottom eqn by 2 and add them together

OpenStudy (dan815):

and u shud get the eqn of a line in i and j

OpenStudy (dan815):

3i + j - 4k = 4 and -2i + 2j - 42k = 4 ------------------- 5i-j=0 j=5i